# A Zeroth-Order Extra-Gradient Method For Black-Box Constrained Optimization

### v1

@@ 1   Introduction
Note: $\tau$ denotes the block size of randomly selected coordinates (see detailed definition in Section 4).
@@ 3.1   Algorithm design
$\mathbf{g}_{k}^{+}=\frac{f(\mathbf{z}_{k}^{+}+r_{k}\mathbf{w}_{k}-f(\mathbf{z}_{k}^{+})}{r_{k}/d}\tilde{\mathbf{w}}_{k},$
@@ 3.2   Finite-sample analysis
Proof of Lemma 3.2.  Following the definition of $(\hat{\mathbf{x}}_{K},\hat{\mathbf{y}}_{K})$, the duality gap $\mathbb{E}\left[f\left(\hat{\mathbf{x}}_{K},\mathbf{y}\right)-f\left(\mathbf{x},\hat{\mathbf{y}}_{K}\right)\right]$ is upper bounded by:
@@
$\|\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]_{+}\|\leq\frac{\Gamma_{K}}{RK}.$
@@ 4.3   Zeroth-order block coordinate extra-gradient
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq$ $\ \left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)\frac{\tilde{D}^{2}}{2\eta K}+2\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)\eta G^{2}+\frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2K}+\frac{d_{x}+d_{y}}{2K}\eta L^{2}M_{2}$
$\ +\frac{2\eta GLM_{1}}{K}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}}).$
Based on the result in Theorem 4.3, ZOBCEG can achieve the convergence to a saddle point at the rate of $\mathcal{O}\left(\frac{\tilde{d}}{\eta K}+\frac{\eta d}{K}+\tilde{d}\eta\right)$, where $\tilde{d}=\max\left\{\frac{d_{x}}{\tau_{x}},\frac{d_{y}}{\tau_{y}}\right\}$. Similarly to ZOEG, we can apply a small step size of $\eta$ to achieve the convergence rate of $\mathcal{O}\left(\frac{\tilde{d}}{\sqrt{K}}\right)$. Furthermore, the oracle complexity of ZOBCEG is $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ to derive an $\epsilon$-saddle point for problem (2) and an $\epsilon$-optimal solution for problem (1).
@@ 6   Conclusions
$\ \mathbb{E}\left[\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (26)
$\leq$ $\ \mathbb{E}\left[\left\langle\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}.$
@@
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}$
$\leq$ $\ \eta\|\tilde{\mathbf{g}}_{k}^{x,+}\|\cdot\|\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x}\|+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}$
$\leq$ $\ \eta\left(2G^{2}+2GLr_{k}\sqrt{\tau_{x}}+\frac{L^{2}r_{k}^{2}\tau_{x}}{2}\right)+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta},$ (27)
where the first inequality follows from the projection property, the second one follows from algebraic transformation, and the last inequality follows from $\|\tilde{\mathbf{g}}_{k}^{x,+}\|\leq G+\frac{Lr_{k}\sqrt{\tau_{x}}}{2}$. Substituting (C) into (LABEL:eq:_block_grad_bound_ZOCEG_term) gives:
$\ \mathbb{E}\left[\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (28)
$\leq$ $\ \frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}+\frac{\eta d_{x}}{\tau_{x}}\left(2G^{2}+2GLr_{k}\sqrt{\tau_{x}}+\frac{L^{2}r_{k}^{2}\tau_{x}}{2}\right).$
@@
$\ \mathbb{E}\left[-\left\langle\nabla_{\mathbf{y}}f(\mathbf{z}_{k}^{+}),\mathbf{y}_{k}^{+}-\mathbf{y}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$ (29)
$\leq$ $\ \frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{k}^{y}(\mathbf{y})-H_{k+1}^{y}(\mathbf{y})\mid\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{y}}\tilde{D}}{2}+\frac{\eta d_{y}}{\tau_{y}}\left(2G^{2}+2GLr_{k}\sqrt{\tau_{y}}+\frac{L^{2}r_{k}^{2}\tau_{y}}{2}\right).$
@@
$\ \mathbb{E}\left[\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\mid\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \frac{Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{k}^{x}(\mathbf{z})-H_{k+1}^{x}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{k}^{y}(\mathbf{z})-H_{k+1}^{y}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]$
$\ +2\eta G^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLr_{k}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})+\frac{d_{x}+d_{y}}{2}\eta L^{2}r_{k}^{2}.$
@@
$\leq$ $\ \frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{0}^{x}(\mathbf{z})-H_{K}^{x}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{0}^{y}(\mathbf{z})-H_{K}^{y}(\mathbf{z})\mid\mathcal{F}_{k}^{b}\right]$
$\ +2\eta KG^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLM_{1}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})+\frac{d_{x}+d_{y}}{2}\eta L^{2}M_{2}$
$\leq$ $\ \frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{d_{x}\tilde{D}^{2}}{2\eta\tau_{x}}+\frac{d_{y}\tilde{D}^{2}}{2\eta\tau_{y}}+2\eta KG^{2}\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)+2\eta GLM_{1}(\frac{d_{x}}{\sqrt{\tau_{x}}}+\frac{d_{y}}{\sqrt{\tau_{y}}})$
$\ +\frac{d_{x}+d_{y}}{2}\eta L^{2}M_{2},$ (30)

### v2

@@ 1   Introduction
Note: $\tau$ denotes the block size (see detailed definition in Section 4).
@@ 3.1   Algorithm Design
$\mathbf{g}_{k}^{+}=\frac{f(\mathbf{z}_{k}^{+}+r_{k}\mathbf{w}_{k})-f(\mathbf{z}_{k}^{+})}{r_{k}/d}\tilde{\mathbf{w}}_{k},$
@@ 3.2   Finite-Sample Analysis
Proof of Lemma 3.2.  Following the definition of $(\hat{\mathbf{x}}_{K},\hat{\mathbf{y}}_{K})$, the duality gap $\mathbb{E}\left[f\left(\hat{\mathbf{x}}_{K},\mathbf{y}^{*}\right)-f\left(\mathbf{x}^{*},\hat{\mathbf{y}}_{K}\right)\right]$ is upper bounded by:
@@
$\|\left[\mathbb{E}\left[\bm{\phi}(\hat{\mathbf{x}}_{K})\right]\right]_{+}\|\leq\frac{\Gamma_{K}}{RK}.$
@@ 4.3   Zeroth-Order Block Coordinate Extra-Gradient
$\mathbb{E}\left[\Delta(\hat{\mathbf{z}}_{K})\right]\leq\left(\frac{d_{x}}{\tau_{x}}+\frac{d_{y}}{\tau_{y}}\right)\frac{\tilde{D}^{2}}{2\eta K}+\frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2K}+\frac{3(d_{x}+d_{y})}{2K}\eta L^{2}M_{2}+24\eta G^{2}.$
Based on the result in Theorem 4.3, ZOBCEG can achieve the convergence to a saddle point at the rate of $\mathcal{O}\left(\frac{\tilde{d}}{\eta K}+\frac{\eta d}{K}+\eta\right)$, where $\tilde{d}=\max\left\{\frac{d_{x}}{\tau_{x}},\frac{d_{y}}{\tau_{y}}\right\}$. Similarly to ZOEG, we can apply a small step size of $\eta$, i.e., $\eta=\mathcal{O}\left(\sqrt{\frac{\tilde{d}}{K}}\right)$, to achieve the convergence rate of $\mathcal{O}\left(\frac{\sqrt{\tilde{d}}}{\sqrt{K}}\right)$. Furthermore, the oracle complexity of ZOBCEG is $\mathcal{O}\left(\frac{d}{\epsilon^{2}}\right)$ to derive an $\epsilon$-saddle point for problem (2) and an $\epsilon$-optimal solution for problem (1).
@@ 6   Conclusions
$\ \mathbb{E}\left[\left.\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\right|\mathcal{F}_{k}^{b}\right]\leq\mathbb{E}\left[\left.\left\langle\frac{d_{x}}{\tau_{x}}\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\right|\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}.$ (26)
@@
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}-\frac{1}{2\eta}\|\mathbf{x}_{k}-\mathbf{x}_{k+1}\|^{2}$
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle-\frac{1}{\eta}\left\langle\mathbf{x}_{k}^{+}-\mathbf{x}_{k},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}-\frac{1}{2\eta}\|\mathbf{x}_{k}-\mathbf{x}_{k+1}\|^{2}$
$\leq$ $\ \left\langle\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x},\mathbf{x}_{k}^{+}-\mathbf{x}_{k+1}\right\rangle+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta}$
$\leq$ $\ \eta\|\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x}\|^{2}+\frac{H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})}{\eta},$ (27)
where the first, third, and last inequalities follow from the projection property, while the second and fourth arise from straightforward algebraic transformations . The conditional expectation of $\|\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x}\|^{2}$ can be bounded as follows:
$\ \mathbb{E}\left[\left.\left\|\tilde{\mathbf{g}}_{k}^{x,+}-\tilde{\mathbf{g}}_{k}^{x}\right\|^{2}\right|\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ 3\cdot\mathbb{E}\left[\left.\left\|\tilde{\mathbf{g}}_{k}^{x,+}-\sum_{p\in\mathcal{J}_{k}^{x}}\nabla_{\mathbf{x},p}f(\mathbf{z}_{k}^{+})\right\|^{2}+\left\|\tilde{\mathbf{g}}_{k}^{x}-\sum_{i\in\mathcal{I}_{k}^{x}}\nabla_{\mathbf{x},i}f(\mathbf{z}_{k})\right\|^{2}\right|\mathcal{F}_{k}^{b}\right]$
$\ +3\cdot\mathbb{E}\left[\left.\left\|\sum_{i\in\mathcal{I}_{k}^{x}}\nabla_{\mathbf{x},i}f(\mathbf{z}_{k})-\sum_{p\in\mathcal{J}_{k}^{x}}\nabla_{\mathbf{x},p}f(\mathbf{z}_{k}^{+})\right\|^{2}\right|\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \frac{3L^{2}r_{k}^{2}\tau_{x}}{2}+6\cdot\mathbb{E}\left[\left.\left\|\sum_{i\in\mathcal{I}_{k}^{x}}\nabla_{\mathbf{x},i}f(\mathbf{z}_{k})\right\|^{2}+\left\|\sum_{p\in\mathcal{J}_{k}^{x}}\nabla_{\mathbf{x},p}f(\mathbf{z}_{k}^{+})\right\|^{2}\right|\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \frac{3L^{2}r_{k}^{2}\tau_{x}}{2}+\frac{6\tau_{x}}{d_{x}}\mathbb{E}\left[\left.\left\|\nabla_{\mathbf{x}}f(\mathbf{z}_{k})\right\|^{2}+\left\|\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+})\right\|^{2}\right|\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \ \frac{3L^{2}r_{k}^{2}\tau_{x}}{2}+\frac{12\tau_{x}G^{2}}{d_{x}},$ (28)
where the second inequality follows from (5). Substituting (C) and (C) into (LABEL:eq:_block_grad_bound_ZOCEG_term) gives:
$\ \mathbb{E}\left[\left.\left\langle\nabla_{\mathbf{x}}f(\mathbf{z}_{k}^{+}),\mathbf{x}_{k}^{+}-\mathbf{x}\right\rangle\right|\mathcal{F}_{k}^{b}\right]$ (29)
$\leq$ $\ \frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[\left.H_{k}^{x}(\mathbf{x})-H_{k+1}^{x}(\mathbf{x})\right|\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{x}}\tilde{D}}{2}+\frac{3\eta L^{2}r_{k}^{2}d_{x}}{2}+12\eta G^{2}.$
@@
$\ \mathbb{E}\left[\left.-\left\langle\nabla_{\mathbf{y}}f(\mathbf{z}_{k}^{+}),\mathbf{y}_{k}^{+}-\mathbf{y}\right\rangle\right|\mathcal{F}_{k}^{b}\right]$ (30)
$\leq$ $\ \frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[\left.H_{k}^{y}(\mathbf{y})-H_{k+1}^{y}(\mathbf{y})\right|\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}\sqrt{d_{y}}\tilde{D}}{2}+\frac{3\eta L^{2}r_{k}^{2}d_{y}}{2}+12\eta G^{2}.$
@@
$\ \mathbb{E}\left[\left.\left\langle F(\mathbf{z}_{k}^{+}),\mathbf{z}_{k}^{+}-\mathbf{z}\right\rangle\right|\mathcal{F}_{k}^{b}\right]$
$\leq$ $\ \frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[\left.H_{k}^{x}(\mathbf{z})-H_{k+1}^{x}(\mathbf{z})\right|\mathcal{F}_{k}^{b}\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[\left.H_{k}^{y}(\mathbf{z})-H_{k+1}^{y}(\mathbf{z})\right|\mathcal{F}_{k}^{b}\right]+\frac{Lr_{k}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}$
$\ +\frac{3(d_{x}+d_{y})}{2}\eta L^{2}r_{k}^{2}+24\eta G^{2}.$
@@
$\leq$ $\ \frac{d_{x}}{\eta\tau_{x}}\mathbb{E}\left[H_{0}^{x}(\mathbf{z})-H_{K}^{x}(\mathbf{z})\right]+\frac{d_{y}}{\eta\tau_{y}}\mathbb{E}\left[H_{0}^{y}(\mathbf{z})-H_{K}^{y}(\mathbf{z})\right]+\frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}$
$\ +\frac{3(d_{x}+d_{y})}{2}\eta L^{2}M_{2}+24\eta KG^{2}$
$\leq$ $\ \frac{d_{x}\tilde{D}^{2}}{2\eta\tau_{x}}+\frac{d_{y}\tilde{D}^{2}}{2\eta\tau_{y}}+\frac{LM_{1}(\sqrt{d_{x}}+\sqrt{d_{y}})\tilde{D}}{2}+\frac{3(d_{x}+d_{y})}{2}\eta L^{2}M_{2}+24\eta KG^{2},$ (31)